### v3

# Reconstructing a pseudotree from the distance matrix of its boundary

Abstract
A vertex $v$ of a connected graph $G$ is said to be a boundary vertex of $G$ if for some other vertex $u$ of $G$, no neighbor of $v$ is further away from $u$ than $v$. The boundary $\partial(G)$ of $G$ is the set of all of its boundary vertices.
The boundary distance matrix $\hat{D}_{G}$ of a graph $G=([n],E)$ is the square matrix of order $\kappa$, being $\kappa$ the order of $\partial(G)$, such that for every $i,j\in\partial(G)$, $[\hat{D}_{G}]_{ij}=d_{G}(i,j)$.
Given a square matrix $\hat{B}$ of order $\kappa$, we prove under which conditions $\hat{B}$ is the distance matrix $\hat{D}_{T}$ of the set of leaves of a tree $T$, which is precisely its boundary.
We show that if $G$ is either a tree or a unicyclic graph with girth $g\geq 5$ vertices, then $G$ is uniquely determined by the boundary distance matrix $\hat{D}_{G}$ of $G$ and we also conjecture that this statement holds for every connected graph $G$, whenever both the order $n$ and the boundary (and thus also the boundary distance matrix) of $G$ are prefixed.
Moreover, an algorithm for reconstructing a tree (resp., a unicyclic graph) from its boundary distance matrix is given, whose time complexity in the worst case is $O(\kappa n)$ (resp., $O(n^{2})$).
1 Introduction
While typically a graph is defined by its lists of vertices and edges, significant research has been dedicated to minimizing the necessary information required to uniquely determine a graph. For example, various approaches include reconstructing metric graphs from density functions [9], road networks from a set of trajectories [1], graphs utilizing shortest paths or distance oracles [15], labeled graphs from all r-neighborhoods [19], or reconstructing phylogenetic trees [2]. Of particular note is the graph reconstruction conjecture [16, 26] which states the possibility of reconstructing any graph on at least three vertices (up to isomorphism) from the multiset of all unlabeled subgraphs obtained through the removal of a single vertex. Indeed, a search with the words “graph reconstruction” returns more than 3 million entries.
In this paper, our focus lies in the reconstruction of graphs from the distance matrix of their boundary vertices and the graph’s order. We are persuaded that this process could hold true for all graphs, and we state it as a conjecture (see Conjecture 11). It is accordingly of particular interest to explore whether this conjecture holds for specific families of graphs. Our objective herein is to establish its validity for pseudotrees, i.e., for both trees and unicyclic graphs.
The concept of a graph’s boundary was introduced by Chartrand, Erwin, Johns and Zhang in 2003 [7]. Initially conceived to identify local maxima of vertex distances, the boundary has since revealed a host of intriguing properties. It has been recognized as geodetic [4], serving as a resolving set [13], and also as a strong resolving set [21]. Put simply, each vertex lies in the shortest path between two boundary vertices and, given any pair of vertices $x$ and $y$, there exists a boundary vertex $v$ such that either $x$ lies on the shortest path between $v$ and $y$, or vice versa. With such properties, it is unsurprising that the boundary emerges as a promising candidate for reconstructing the entire graph.
Graph distance matrices represent a fundamental tool for graph users, enabling the solution of problems such as finding the shortest path between two nodes. However, our focus here shifts mainly towards their realizability. That is, given a matrix (integer, positive and symmetric), we inquire whether there exists a corresponding graph where the matrix entries represent the distances between vertices. In 1965, Hakimi and Yau [11] presented a straightforward additional condition that the matrix must satisfy to be realizable (see Theorem 7). Building upon this, in 1974, Buneman [3] provided the matrix characterization for being the distance matrix of a tree once we know that the graph is $K_{3}$-free, and Graham and Pollack [10] computed the determinant of the distance matrix of a tree (see Theorem 16). Additionally, Howorka [14] in 1979, formulated conditions for the distance matrix of a block graph (see Theorem 14), and Lin, Liu and Lu [18] provided the determinant of such matrices (see Theorem 15). Incidentally, we use their result to derive the converse of the Graham and Pollack’s theorem. Furthermore, we also give an algorithmic approach to the characterization of the distance matrix of a unicyclic graph.
As previously mentioned, we are interested in the distance matrix of a graph’s boundary, a submatrix of the overall distance matrix. We seek to determine the realizability of these matrices and, while we have achieved characterization in the case of trees, a similar analysis for unicyclic graphs remains elusive.
Finally, we present a pair of algorithms for reconstructing trees and unicyclic graphs from the distance matrix of their boundary. In trees, the boundary corresponds to the leaves, while in unicyclic graphs, it contains the leaves along with the vertices of the cycle with degree two. Notably, the concept of doubly resolving sets, introduced in [5], proves instrumental in these reconstructions.
The paper is organized as follows: this section is finished by introducing general terminology and notation. In Section 2, we explore the notion of boundary and its relation with distance matrices. Section 3 is devoted to the reconstruction of trees, completing first with the realizability of both the distance matrix and the boundary distance matrix of a tree. In a similar way, Section 4 undertakes the characterization of distance matrices for unicyclic graphs, followed by their reconstruction from the boundary distance matrix. Finally, the paper ends with a section on conclusions and open problems.
1.1 Basic terminology
All the graphs considered are undirected, simple, finite and (unless otherwise stated) connected. If $G=(V,E)$ is a graph of order $n$ and size $m$, it means that $|V|=n$ and $|E|=m$. Unless otherwise specified, $n\geq 2$ and $V=[n]=\{1,\ldots,n\}$.
Let $v$ be a vertex of a graph $G$. The open neighborhood of $v$ is $N(v)=\{w\in V(G):vw\in E\}$, and the closed neighborhood of $v$ is $N[v]=N(v)\cup\{v\}$ The degree of $v$ is $\deg(v)=|N(v)|$. The minimum degree (resp. maximum degree) of $G$ is $\delta(G)=\min\{\deg(u):u\in V(G)\}$ (resp. $\Delta(G)=\max\{\deg(u):u\in V(G)\}$). If $\deg(v)=1$, then $v$ is said to be a leaf of $G$ and the set and the number of leaves of $G$ are denoted by ${\cal L}(G)$ and $\ell(G)$, respectively.
Let $K_{n}$, $P_{n}$, $W_{n}$ and $C_{n}$ be, respectively, the complete graph, path, wheel and cycle of order $n$. Moreover, $K_{r,s}$ denotes the complete bipartite graph whose maximal independent sets are $\overline{K_{r}}$ and $\overline{K_{s}}$, respectively. In particular, $K_{1,n-1}$ denotes the star with $n-1$ leaves.
Given a graph $G=(V,E)$ and a subset of vertices $W\subseteq V$, the subgraph of $G$ induced by $W$, denoted by $G[W]$, has $W$ as vertex set and $E(G[W])=\{vw\in E:v,w\in W\}$. If $G[W]$ is a complete graph, then it is said to be a clique of $G$.
Given a pair of vertices $u,v$ of a graph $G$, a $u-v$ geodesic is a $u-v$ shortest path, i.e., a path joining $u$ and $v$ of minimum order. Clearly, all $u-v$ geodesics have the same length, and it is called the distance between vertices $u$ and $v$ in $G$, denoted by $d_{G}(u,v)$, or simply by $d(u,v)$, when the context is clear. A set $W\subseteq V(G)$ is called geodetic if any vertex of the graph is in a $u-v$ geodesic for some $u,v\in W$.
The eccentricity $\rm ecc(v)$ of a vertex $v$ is the distance to a farthest vertex from $v$. The radius and diameter of $G$ are respectively, the minimum and maximum eccentricity of its vertices and are denoted as ${\rm rad}(G)$ and ${\rm diam}(G)$. A vertex $u\in V(G)$ is a central vertex of $G$ if $\rm ecc(u)={\rm rad}(G)$, and it is called a peripheral vertex of $G$ if $\rm ecc(u)={\rm diam}(G)$. The set of central (resp., peripheral) vertices of $G$ is called the center (resp., periphery) of $G$.
Let $S=\{w_{1},w_{2},\ldots,w_{k}\}$ be a set of vertices of a graph $G$. The distance $d(v,S)$ between a vertex $v\in V(G)$ and $S$, is the minimum of the distances between $v$ and the vertices of $S$, that is, $d(v,S)=\min\{d(v,w):w\in S\}$. The metric representation $r(v|S)$ of a vertex $v$ with respect to $S$ is defined as the $k$-vector $r(v|S)=(d_{G}(v,w_{1}),d_{G}(v,w_{2})\ldots,d_{G}(v,w_{k}))$. A set of vertices $S$ of $G$ is called resolving if for every pair of distinct vertices $x,y\in V(G)$, $r(x|S)\neq r(y|S)$.
Resolving sets were first introduced by Slater in [24], and since then, many other similar concepts have been defined, particularly, we deal with doubly resolving [5] and strong resolving sets [22, 20] in this paper. A set $S$ is called doubly resolving if for every pair $x,y\in V(G)$, there exist $u,v\in S$ such that $d(x,u)-d(y,u)\neq d(x,v)-d(y,v)$, or equivalently, if $r(x|S)-r(y|S)\neq(a,\stackrel{{\scriptstyle|S|}}{{\ldots}},a)$, for some integer $a$. On the other hand, $S$ is called strong resolving if for every pair $x,y\in V(G)$ either $x$ is in a $y-v$ geodesic or $y$ is in $x-v$ geodesic, for some vertex $v\in S$. In other words, $S$ is a strong resolving set if, for every pair $x,y\in V(G)$ , there is a vertex $v\in S$, such that either $d(x,v)=d(x,y)+d(y,v)$ or $d(y,v)=d(y,x)+d(x,v)$. Clearly, every strong (resp., doubly resolving) set is also resolving, but the converse is far from being true (see Figure 1).
A graph $G$ whose order and size are equal is called unicyclic. These graphs contain a unique cycle that is denoted as $C_{g}$, where $g$ is the girth of $G$. The connected component of $G-E(C_{g})$ containing a vertex $v\in V(C_{g})$ is denoted as $T_{v}$ and it is called the branching tree of $v$. The tree $T_{v}$ is said to be trivial if $V(T_{v})=\{v\}$. A vertex $v\in V(G)$ is a branching vertex if either $v\not\in V(C_{g})$ and $\deg(v)\geq 3$ or $v\in V(C_{g})$ and $\deg(v)\geq 4$.
A graph $G$ of order $n$ and size $m$ is called a pseudotree if $m-1\leq n\leq m$, i.e., if it is either a tree (including a path) or a unicyclic graph (including a cycle).
A cut-vertex is a vertex whose deletion disconnects the graph. If $G$ has no cut-vertices, then it is called a block. A maximal subgraph of $G$ without cut-vertices is a block of $G$. In a block graph, every block is a clique, or equivalently, every cycle induces a complete subgraph.
For further information on basic Graph Theory we refer the reader to [8].
2 The conjecture
2.1 Boundary of a graph
In this subsection, we introduce one of the essential components of our work: the boundary of a graph, which was first studied by Chartrand et al. in [7]. A vertex $v$ of a graph $G$ is said to be a boundary vertex of a vertex $u$ if no neighbor of $v$ is further away from $u$ than $v$, i.e., if for every vertex $w\in N(v)$, $d(u,w)\leq d(u,v)$. The set of boundary vertices of a vertex $u$ is denoted by $\partial(u)$. Given a pair of vertices $u,v\in V(G)$ if $v\in\partial(u)$, then $v$ is also said to be maximally distant from $u$. A pair of vertices $u,v\in V(G)$ are called mutually maximally distant, or simply $MMD$, if both $v\in\partial(u)$ and $u\in\partial(v)$.
The boundary of $G$, denoted by $\partial(G)$, is the set of all of its boundary vertices, i.e., $\partial(G)=\cup_{u\in V(G)}\partial(u)$. Notice that, as was pointed out in [21], the boundary of $G$ can also be defined as the set of $MMD$ vertices of $G$, i.e.,
$\partial(G)=\{v\in V(G):{\rm there\,exists}\,u\in V(G)\,{\rm such\,that}\,u,v\,{\,\rm are}\,MMD\}$
Theorem 1 ([12, 25]).
Let $G$ be a graph of order $n\geq 2$ with $\kappa$ boundary vertices. Then, $\kappa=2$ if and only if $G=P_{n}$. Moreover, $\kappa=3$ if and only if either
  1. 1.
    $G$ is a subdivision of $K_{1,3}$; or
  2. 2.
    $G$ can be obtained from $K_{3}$ by attaching exactly one path (of arbitrary length) to each of its vertices.
Also, graphs with a big $\kappa$ are well-known, as the next results show.
Proposition 2.
Given a graph $G$ of order $n$ with $\kappa$ boundary vertices,
  1. 1.
    If ${\rm rad}(G)={\rm diam}(G)$, then $\kappa=n$.
  2. 2.
    If ${\rm diam}(G)=2$, then $n-1\leq\kappa\leq n$.
    Moreover, $\kappa=n-1$ if and only if contains a unique central vertex.
Proof.
  1. 1.
    Suppose that ${\rm rad}(G)={\rm diam}(G)=d$. Take $u\in V(G)$ and notice that $\rm ecc(u)=d$. Let $v\in V(G)$ such that $d(u,v)=d$. For every vertex $w\in N(v)$, $d(u,w)\leq d=d(u,v)$. Hence, $u\in\partial(u)\subseteq\partial(G)$.
  2. 2.
    If ${\rm rad}(G)={\rm diam}(G)=2$, then according to the previous item, $\kappa=n$. Suppose that ${\rm rad}(G)=1$. Let $V(G)=U\cup W$ such that $U$ is the set of central vertices and $W$ is the set of peripheral vertices of $G$. Observe that $W\subsetneq\partial(G)$ and that if $|U|=h$, then $G[U)]=K_{h}$. If $h\geq 2$, then every central vertex belongs to the boundary of every other vertex of $G$. If $h=1$ and $U=\{u\}$, then for every vertex $w\in W$, $u\not\in\partial(w)$, i.e., $\partial(G)=W$, which means that $\kappa=|W|=n-1$.
Corollary 3.
Let $G$ be a graph of order $n\geq 3$ with $\kappa$ boundary vertices.
  1. 1.
    If $G\in\{K_{n},K_{r,s},C_{n}\}$ and $r,s\geq 2$, then $\kappa=n$.
  2. 2.
    If $G\in\{W_{n},K_{1,n-1}\}$, then $\kappa=n-1$.
As previously mentioned, the boundary exhibits several intriguing properties, like being geodetic [4] and a resolving set [13]. However, for the scope of this paper, its status as a strong resolving set is particularly pertinent. Thus, we shall now develop into this concept with some detail.
That notion were first defined by Sebő and Tannier [22] in 2003, and later studied in [20]. They were interested in extending isometric embeddings of subgraphs into the whole graph and, to ensure that, they defined a strong resolving set of a graph $G$ as a subset $S\subseteq V(G)$ such that for any pair $x,y\in V(G)$ there is an element $v\in S$ such that there exists a $x-v$ geodesic that contains $y$, or a $y-v$ geodesic containing $x$.
What is crucial for our goals is that, as a consequence of the definition, it only suffices to know the distances from the vertices of a strong resolving set to the rest of nodes, to uniquely determine the graph. This issue is explored in more detail in Subsection 2.2.
It turns out that the boundary of a graph is always a strong resolving set.
Proposition 4 ([21]).
The boundary $\partial(G)$ of every graph $G$ is a strong resolving set.
Proof.
Let $u,v\in V(G)$ such that $d(u,v)=k$ and $\{u,v\}\cap\partial(G)=\emptyset$. So, for some vertex $w_{1}\in N(v)$, $d(u,w_{1})=k+1$. If $w_{1}\in\partial(u)$, then we are done. Otherwise, for some vertex $w_{2}\in N(w_{1})$, $d(u,w_{2})=k+2$.
Thus, after iterating this procedure finitely many times, say $h$ times, we will finally find a vertex $w_{h}$ such that for every vertex $w\in N(w_{h})$, $d(u,w)\leq d(u,w_{h})=k+h$, i.e., a vertex $w_{h}\in\partial(G)$ and a $u-w_{h}$ geodesic containing vertex $v$. ∎
Particularly, for trees and unicyclic graphs, the boundary is very straightforward to characterize.
Proposition 5 ([25]).
Let $T$ be a tree. Then, $\partial(T)={\cal L}(T)$.
Proof.
If $u\in{\cal L}(T)$ and $N(u)=\{v\}$, then notice that $u\in\partial(v)$, and thus $u\in\partial(T)$.
Take a vertex $u\in V(T)$ such that $\deg(u)\geq 2$. If $\{v_{1},v_{2}\}\subseteq N(u)$ then, for every vertex $w\in V(T)$, $d(w,u)<\max\{d(w,v_{1}),d(w,v_{2})\}$. Hence, $u\not\in\partial(G)$. ∎
Proposition 6.
Let $G$ be a unicyclic graph of girth $g$.. If ${\cal U}(G)$ denotes the set of vertices of $C_{g}$ of degree 2, then $\partial(G)={\cal L}(G)\cup{\cal U}(G)$.
Proof.
If $u\in{\cal L}(G)$ and $N(u)=\{v\}$, then notice that $u\in\partial(v)$.
Let $u\in{\cal U}(G)$. Let $v\in V(C_{g})$ such that $d(u,v)\in\lfloor\frac{g}{2}\rfloor$. If $w$ is a leaf of the branching tree $T_{v}$ of $v$, then observe that $u\in\partial(w)$.
Finally, take a vertex $u\not\in{\cal L}(G)\cup{\cal U}(G)$. If $u\in V(C_{g})$, $N(u)\cap V(C_{g})=\{v_{1},v_{2}\}$ and $v_{3}\in N(u)\cap V(T_{u})$, then $d(w,u)<\max\{d(w,v_{1}),d(w,v_{2}),d(w,v_{3})\}$, for every vertex $w\in V(G)$. Thus, $u\not\in\partial(G)$. If $u\in V(T_{v})$ for some vertex $v\in V(C_{g})$, $\deg(u)\geq 2$ and $\{v_{1},v_{2}\}\subseteq N(u)$, then $d(w,u)<\max\{d(w,u_{1}),d(w,v_{2})\}$, for every vertex $w\in V(G)$. Hence, $u\not\in\partial(G)$. ∎
2.2 The distance matrix of a graph
At this point, the other relevant element of the work is introduced: distance matrices. Some notation is provided as well as a complete characterization of both the distance matrix of a tree and the distance matrix of the leaves of a tree. This subsection concludes by showing a conjecture, along with some related open problems.
A square matrix $D$ is called a dissimilarity matrix if it is symmetric, all off-diagonal entries are (strictly) positive and the diagonal entries are zeroes. A square matrix $D$ of order $n$ is called a metric dissimilarity matrix if it satisfies, for any triplet $i,j,k\in[n]$, the triangle inequality: $d_{ik}\leq d_{ij}+d_{jk}$.
The distance matrix $D_{G}$ of a graph $G=(V,E)$ with $V=[n]$ is the square matrix of order $n$ such that, for every $i,j\in[n]$, $d_{ij}=d(i,j)$. Certainly, this matrix is a metric dissimilarity matrix. A metric dissimilarity matrix $D$ is called a distance matrix if there is graph $G$ such that $D_{G}=D$.
Let $S$ be a subset of vertices of order $k$ of a graph $G=(V,E)$, with $V=[n]$. It is denoted by $D_{S,V}$ the submatrix of $D_{G}$ of order $k\times n$ such that for every $i\in S$ and for every $j\in V$, $[D_{S,V}]_{ij}=d(i,j)$.
Similarly, the so-called $S$-distance matrix of $G$, denoted by $D^{G}_{S}$, or simply by $D_{S}$, when the context is clear, is the square submatrix of $D_{G}$ of order $k$ such that for every $i,j\in S$, $[D_{S}]_{ij}=d(i,j)$. If $S=\partial(G)$, then $D_{S}$ is also denoted by $\hat{D}_{G}$ and it is called the boundary distance matrix of $G$.
The next result was stated and proved in [11] and constitutes a general characterization of distance matrices. We include here a (new) proof, for the sake of completeness.
Theorem 7 ([11]).
Let $D$ be an integer metric dissimilarity matrix of order $n$. Then, $D$ is a distance matrix if and only if, for every $i,j\in[n]$, if $d_{ij}>1$, then there exists an integer $k\in[n]$ such that
$d_{ik}=1$ and $d_{ij}=d_{ik}+d_{kj}$.
Proof.
The necessity of the above condition immediately follows from the definition of distance matrix.
To prove the sufficiency, we consider the non-negative symmetric square matrix $A$ of order $n$, such that, for every pair $i,j\in[n]$, $a_{ii}\in\{0,1\}$ being $a_{ij}=1$ if and only if $d_{ij}=1$. Let $G$ the graph $G=([n],E)$ of order $n$ such that its adjacency matrix is $A$. Next, we show that the distance matrix of $G$ is precisely $D$.
If ${\rm diam}(G)=d$, then $d(i,j)=p\in\{1,\ldots,d\}$, for every pair of distinct vertices $i,j\in[n]$. If $p=1$, then clearly $d(i,j)=1$ if and only if $d_{ij}=1$. Take $2\leq p\leq d$ and suppose that, if $11\leq r\leq p-$ 1, then
$d(i,j)=r\,\,{\rm if\,\,and\,\,only\,\,if}\,\,d_{ij}=r$
Let $i,j\in[n]$ such that $d_{ij}=p$. According to item (c), take $k\in[n]$ such that $d(i,k)=d_{ik}=1$ and $p=d_{ij}=d_{ik}+d_{kj}=1+d_{kj}$. This means that $d(i,j)\leq p$, since $d_{kj}=p-1$ and $d(i,j)\leq d(i,k)+d(k,j)=1+d(k,j)$. Hence, $d(i,j)=p$ as otherwise, according to the inductive hypothesis (1), $d_{ij}=d(i,j)<p$, a contradiction.
Conversely, let $i,j\in[n]$ such that $d(i,j)=p$. Let $k\in N(i)$ such that $d(i,j)=1+d(k,j)$. According to item (b), $d_{ij}\leq d_{ik}+d_{kj}$. This means that $d_{ij}\leq p$, since $d(k,j)=p-1$ and $d_{ij}\leq 1+d_{kj}$. Hence, $d_{ij}=p$ as otherwise, according to the inductive hypothesis (1), $d(i,j)=d_{ij}<p$, a contradiction. ∎
An integer metric dissimilarity matrix $D$ of order $n$ is called additive if every subset of indices $\{i,j,h,k\}\subseteq[n]$ satisfies the so-named four-point condition:
$\left\{\begin{array}[]{c}d_{ij}+d_{hk}\leq\max\{d_{ih}+d_{jk},d_{ik}+d_{jh}\}\\ \\ d_{ih}+d_{jk}\leq\max\{d_{ij}+d_{hk},d_{ik}+d_{jh}\}\\ \\ d_{ik}+d_{jh}\leq\max\{d_{ij}+d_{hk},d_{ih}+d_{jk}\}\\ \end{array}\right.$
A graph $G=([n],E)$ is said to satisfy the four-point condition if its distance matrix $D_{G}$ is additive, that is, if every 4-vertex set $\{i,j,h,k\}\subseteq[n]$ satisfies this condition, a strengthened version of the triangle inequality (see [3]):
$\left\{\begin{array}[]{c}d(i,j)+d(h,k)\leq\max\{d(i,h)+d(j,k),d(i,k)+d(j,h)\}\\ \\ d(i,h)+d(j,k)\leq\max\{d(i,j)+d(h,k),d(i,k)+d(j,h)\}\\ \\ d(i,k)+d(j,h)\leq\max\{d(i,j)+d(h,k),d(i,h)+d(j,k)\}\\ \end{array}\right.$
As was pointed out in [3], these inequalities can be characterized as follows.
Proposition 8 ([3]).
Let $\{i,j,h,k\}$ be a 4-vertex set of a graph $G=([n],E)$. Then, the following statements are equivalent.
  1. 1.
    $\{i,j,h,k\}$ satisfies the four-point condition.
  2. 2.
    Among the three sums $d(i,j)+d(h,k)$, $d(i,h)+d(j,k)$, $d(i,k)+d(j,h)$, the two largest ones are equal.
The next result was implicitly mentioned in some papers [17, 21, 22] and proved in [6]. This equivalence, along with the statement shown in Proposition 4, has served as an inspiration for the main conjecture of the paper that is presented at the end of this subsection.
Theorem 9 ([6]).
Let $S$ be a proper subset of vertices of a graph $G=(V,E)$. Then, the following statements are equivalent.
  1. 1.
    $S$ is a strong resolving set.
  2. 2.
    $G$ is uniquely determined by the distance matrix $D_{S,V}$.
As was noticed in [21, 22], this result is not true if we consider resolving sets instead of strong resolving sets. For example, the pair of leaves of the graphs displayed in Figure 1 form a resolving set $S$ in both cases, and also for both graphs the matrix $D_{S,V}$ is the same.
Figure 1: A pair of graphs of order $7$, whose pair of leaves form a minimum (not strong) resolving set.
As a direct consequence of both Theorem 9 and Proposition 4, the following result holds.
Corollary 10 ([6]).
Let $G=(V,E)$ be a graph. Then, $G$ is uniquely determined by the distance matrix $D_{\partial(G),V}$.
Figure 2: In each column, the upper and middle graphs have the same boundary but different order, meanwhile the middle and lower graphs have the same order but different boundary. In all cases, the boundary is the set of black vertices.
It is relatively easy two find pairs of graphs having the same boundary (that is, the same boundary distance matrix) but different order (see Figure 2, for some examples). Having in mind all of these results and particularly the one stated in Corollary 10, we present the following conjecture.
Conjecture 11.
Let $\hat{B}$ an integer metric dissimilarity matrix of order $\kappa$. Let $G=([n],E)$ be a graph such that $\hat{D}_{G}=\hat{B}$. If $G^{\prime}=([n],E^{\prime})$ is a graph such that $\hat{D}_{G^{\prime}}=\hat{B}$, then $G$ and $G^{\prime}$ are isomorphic.
Let $\kappa,n$ be integers such that $2\leq\kappa\leq n$. Let $G=([n],E)$ be a graph of order $n$ and $W$ a subset of $[n]$. Let $\hat{B}$ be an integer metric dissimilarity matrix of order $\kappa$. We define the following graph families, denoted by ${\cal H}_{1}$ and ${\cal H}_{2}$, respectively.
  • $G\in{\cal H}_{1}$ if it is the unique graph (up to isomorphism) of order $n$ such that $\hat{D}_{G}=\hat{B}$.
  • $G\in{\cal H}_{2}$ if it is the unique graph (up to isomorphism) of order $n$ such that $\partial(G)=W$ and $\hat{D}_{G}=\hat{B}$.
Notice that Conjecture 11 can be restated as follows.
Conjecture 11. Every graph belongs to ${\cal H}_{2}$.
Although it is not difficult to find graphs not belonging to the graph family ${\cal H}_{1}$ (see Figure 2, for some examples), we are persuaded that, for a wide spectrum of graph classes, it is possible to obtain the whole graph $G$ from its boundary distance matrix. In Sections 3 and 4, we prove that both the tree and the unicyclic families belong to ${\cal H}_{1}$.
3 Realizability and reconstruction of trees
This section is divided into three subsections: in the first one, we revise the main results regarding the characterization of tree distance matrices and we give the converse of the result of Graham and Pollack [10] in Theorem 18. The next subsection is devoted to determine those matrices which can be the ${\cal L}(T)$-distance matrix of a tree $T$, which are given in Theorem 22. Finally, in the last subsection, we describe and check the validity of the procedure to reconstruct a tree having its ${\cal L}(T)$-distance matrix as the only information. That completes our study on trees.
3.1 The distance matrix of a tree
In the seminal paper [3], Peter Buneman noticed that trees satisfy the four-point condition and also showed that a $K_{3}$-free graph is a tree if and only its distance matrix is additive. In the same paper, it was also proved that, for every additive matrix $A$ of order $k$, there always exists a weighted tree of order $n\geq k$ containing a subset of vertices $S$ of order $k$ such that $D_{S}=A$ (see Figure 3). A different approach based on the structure of the $4\times 4$ principal submatrices was given by Simões Pereira in [23]. In addition, it was proved in [27] that for every dissimilarity matrix $D$, it satisfies de four-point condition if and only if there is a unique weighted binary tree $T$ whose $\partial(T)$-distance matrix is $D$.
Starting from these results, Edward Howorka in [14] was able to characterize the family of graphs whose distance matrix is additive, i.e., satisfying the four-point condition. We include next new proofs of those results for the sake of both completeness and clarity.
Figure 3: A pair of graphs $G_{1}$ and $G_{2}$ of order 4 and 5 respectively, being $G_{2}$ a weighted tree, such that $D_{G_{1}}$, the distance matrix of $G_{1}$, and $D_{[4]}$, the $[4]$-distance matrix of $G_{2}$, are the same.
Proposition 12 ([14]).
Every block graph satisfies the four-point condition.
Proof.
Let $S$ be a 4-vertex set of a block graph $G$, named $S=\{1,2,3,4\}$. The only seven possible configurations of paths connecting the 4 vertices of $S$ are those shown in Figure 4. We check that the four-condition holds in all cases.
  1. 1.
    $d_{12}+d_{34}=d_{13}+d_{24}=d_{14}+d_{23}=a+b+c+d$
  2. 2.
    $d_{12}+d_{34}=a+b+c+d$
    $d_{13}+d_{24}=d_{14}+d_{23}=a+b+c+d+2e$
  3. 3.
    $d_{12}+d_{34}=a+b+c+d$
    $d_{13}+d_{24}=d_{14}+d_{23}=a+b+c+d+2$
  4. 4.
    $d_{12}+d_{34}=a+b+c+d+1$
    $d_{13}+d_{24}=d_{14}+d_{23}=a+b+c+d+2e+2$
  5. 5.
    $d_{12}+d_{34}=a+b+c+d$
    $d_{13}+d_{24}=d_{14}+d_{23}=a+b+c+d+2$
  6. 6.
    $d_{12}+d_{34}=a+b+c+d+2$
    $d_{13}+d_{24}=d_{14}+d_{23}=a+b+c+d+2e+4$
  7. 7.
    $d_{12}+d_{34}=d_{13}+d_{24}=d_{14}+d_{23}=a+b+c+d+2$
Figure 4: There are seven possible configurations of paths connecting 4 vertices of a block graph.
The converse is proved in the next proposition.
Proposition 13 ([14]).
If $G$ satisfies the four-point condition, then it is a block graph.
Proof.
Let $C_{h}$ be an induced cycle of $G$ of minimum order $h\geq 4$. Then, $h=4q+r$, with $q\geq 1$ and $0\leq r\leq 3$. Notice that $C_{h}$ is not only an induced subgraph of $G$ but also isometric. Take a 4-vertex set $\{i,j,h,k\}\subseteq V(C_{h})$ such that $\{d_{ij},d_{jh},d_{h,k},d_{k,i}\}\subseteq\{q,q+1\}$. Check that $d_{ij}+d_{hk}\leq 2q+2$, $d_{ik}+d_{jh}\leq 2q+2$ and $d_{ih}+d_{jk}\geq 4q$. Hence, this 4-vertex set violate the four-point condition, which means that either $G$ is a tree or it is a chordal graph of girth 3, i.e., the only induced cycles have length 3.
Next, suppose that $G$ is a chordal graph of girth 3. Take a cycle $C_{p}=([p],E)$ in $G$ of minimum order $p\geq 4$, such that $[p]$ is not a clique. Notice that $p\geq 5$, since neither the cycle $C_{4}$ nor the diamond $K_{4}-e$ satisfies the four-point condition. Let $i,j\in[p]$ such that $1\leq i<j\leq p$ and $ij\not\in E(G)$. Notice that $d(i,j)=2$, since $C_{p}$ is of minimum order. W.l.o.g. we may assume that $i=1$ and $j=3$. Observe that for every $h\not\in\{2,p\}$ and $k\not\in\{2,4\}$, $\{1h,3k\}\cap E(G)=\emptyset$, since $C_{p}$ is of minimum order.
Let $h$ be the minimum integer between $4$ and $p$ such that $2h\in E(G)$. Then, clearly $h=4$, since otherwise the set $\{2,\ldots,h\}$ is an induced cycle of order at least 4, a contradiction. Let $k$ be the minimum integer between $5$ and $p$ such that $2k\in E(G)$. We distinguish cases.
Case 1. If $k=5$, then the subgraph induced by the set $\{2,3,4,5\}$ is the diamond $K_{4}-e$, a contradiction.
Case 2. If $p=5$ and $2k\in E(G)$, then the subgraph induced by the set $\{2,4,5,1\}$ is the cycle $C_{4}$, a contradiction.
Case 3. If $p\geq 6$ and $k\geq 6$, then the subgraph induced by the set $\{2,4,5,\ldots,k\}$ is the cycle $C_{k-2}$, a contradiction.
Hence, we have proved that every cycle of $G$ induces a clique, i.e., $G$ is a block graph. ∎
Once the two implications have been proved, we can establish the theorem.
Theorem 14 ([14]).
A graph $G$ of order $n$ is a block graph if and only if its distance matrix $D_{G}$ is additive.
Theorem 15 ([18]).
Let $G$ be a block graph on $n$ vertices and $k$ blocks $K_{n_{1}},\ldots,K_{n_{k}}$. Then,
$\det({D_{G}})=(-1)^{n-1}\sum_{i=1}^{k}\frac{n_{i}-1}{n_{i}}\prod_{j=1}^{k}n_{j}$
In particular, as a straight consequence of the previous result, the following theorem, proved in [10], is obtained.
Theorem 16 ([10]).
Let $T$ be a tree on $n$ vertices.Then,
$\det({D_{T}})=(-1)^{n-1}(n-1)2^{n-2}$
The next lemma is the crucial result that allows us to prove the characterization of distance matrices of trees by means of its determinant.
Lemma 17.
Let $k$ and $n$ integers such that $n\geq 3$ and $11\leq k\leq n-$ 1. Let $\{n_{1},\ldots,n_{k}\}$ a decreasing sequence of $k$ integers such that $n-1\geq n_{1}\geq\ldots\geq n_{k}\geq 2$ and $n_{1}+\ldots+n_{k}=n+k-1$. Then,
$\sum_{i=1}^{k}\frac{n_{i}-1}{n_{i}}\prod_{j=1}^{k}n_{j}\leq(n-1)2^{n-2}$
Moreover, the equality holds if and only if $k=n-1$ and $n_{1}=\ldots=n_{n-1}=2$.
Proof.
Let $h\in[k]$ such that $n_{h}\geq 3$ and for every $i\in\{h+1,\ldots,k\}$, $n_{i}=2$. Then,
$\sum_{i=1}^{k}\frac{n_{i}-1}{n_{i}}\prod_{j=1}^{k}n_{j}=\Big{(}\frac{n_{1}-1}{n_{1}}+\ldots+\frac{n_{h}-1}{n_{h}}+\frac{k-h}{2}\Big{)}\cdot\prod_{j=1}^{h}n_{j}\cdot 2^{k-h}$
Take the $(k+1)$-sequence $\{n^{\prime}_{1},\ldots,n^{\prime}_{k+1}\}=\{n_{1},\ldots,n_{h-1},n_{h}-1,n_{h+1},\ldots,n_{k},2\}$. Then,
$\sum_{i=1}^{k+1}\frac{n^{\prime}_{i}-1}{n^{\prime}_{i}}\prod_{j=1}^{k+1}n^{\prime}_{j}=\Big{(}\frac{n_{1}-1}{n_{1}}+\ldots+\frac{n_{h-1}-1}{n_{h-1}}+\frac{n_{h}-2}{n_{h}-1}+\frac{k-h}{2}+\frac{1}{2}\Big{)}\cdot\prod_{j=1}^{h-1}n_{j}\cdot\frac{n_{h}-2}{n_{h}-1}\cdot 2^{k-h+1}$
Check that if $n_{h}\geq 3$, then both $\frac{n_{h}-1}{n_{h}}<\frac{n_{h}-2}{n_{h}-1}+\frac{1}{2}$ and $n_{h}<(n_{h}-1)\cdot 2$.
Hence, $\sum_{i=1}^{k}\frac{n_{i}-1}{n_{i}}\prod_{j=1}^{k}n_{j}<\sum_{i=1}^{k+1}\frac{n^{\prime}_{i}-1}{n^{\prime}_{i}}\prod_{j=1}^{k+1}n^{\prime}_{j}$.
Repeating this procedure iteratively, starting from the sequence $\{n^{\prime}_{1},\ldots,n^{\prime}_{k+1}\}$, the inequality $\sum_{i=1}^{k}\frac{n_{i}-1}{n_{i}}\prod_{j=1}^{k}n_{j}\leq(n-1)2^{n-2}$ is shown, since the last sequence is the $(n-1)$-sequence: $\{2,\ldots,2\}$. ∎
As a direct consequence of Theorems 14, 15, 16 and Lemma 17, we are able to prove the converse of Theorem 16.
Theorem 18.
A graph of order $n$ is a tree $T$ if and only if its distance matrix $D_{T}$ is additive and $\det({D_{T}})=(-1)^{n-1}(n-1)2^{n-2}$.
3.2 Boundary distance matrix of a tree
If, in the previous subsection, we achieved a characterization of the distance matrix of a tree $T$, in this one we intend to characterize the set of metric dissimilarity matrices which are the distance matrix of the set of leaves of a tree. Let us start by showing that no two (non-isomorphic) trees can have the same boundary distance matrix.
Theorem 19.
Let $T$ be a tree on $n$ vertices and $\kappa$ leaves. Then, $T$ is uniquely determined by $\hat{D}_{T}$, the ${\cal L}(T)$-distance matrix of $T$.
Proof.
We proceed by induction on ${\kappa}$. Clearly, the claim holds true when $\kappa=2$ since the unique tree with 2 leaves of order $n$ is the path $P_{n}$ and $n$ is uniquely determined by the distance between its leaves.
Let $T_{\kappa}$ be a tree with $\kappa$ leaves such that ${\cal L}(T_{\kappa})=\{\ell_{1},\ldots,\ell_{\kappa}\}$ is the set of leaves of $T_{\kappa}$. Assume that $\hat{D}_{T_{\kappa}}=\hat{D}_{T}$. Let $\hat{D}_{{\kappa}-1}$ be the submatrix of $\hat{D}_{T}$ obtained by deleting the last row and column of $\hat{D}_{T}$.
By the inductive hypothesis, there is a unique tree $T_{{\kappa}-1}$ with ${\kappa}-1$ leaves such that $\hat{D}_{T_{{\kappa}-1}}=\hat{D}_{{\kappa}-1}$. Hence, $T_{{\kappa}-1}$ is the subtree of $T$ obtained by deleting the path that joins the leaf $\ell_{\kappa}$ to its exterior major vertex $w_{\kappa}$.
According to Proposition 5 and 4, ${\cal L}(T_{\kappa-1})=\{\ell_{1},\ldots,\ell_{\kappa-1}\}$ is a resolving set of $T_{\kappa-1}$. This means that, if $d(\ell_{\kappa},w_{\kappa})=a$, then $w_{\kappa}$ is the unique vertex in $T_{\kappa-1}$ such that
$r(\ell_{\kappa}|{\cal L}(T_{\kappa-1}))=r(w_{\kappa}|{\cal L}(T_{\kappa-1}))+(a,\stackrel{{\scriptstyle\kappa-1}}{{\ldots}},a)$
Thus, $T_{\kappa}$ and $T$ are isomorphic. ∎
Next, we aim to characterize the set of metric dissimilarity matrices which are the distance matrix of the set of leaves of a tree. In order to do that, we first solve the simplest cases, that is, for matrices of order either 3 or 4.
Lemma 20.
Let $\hat{B}$ be an integer metric dissimilarity matrix of order 3. Then, $\hat{B}$ is $\hat{D}_{T}$, the ${\cal L}(T)$-distance matrix of a tree $T$, if and only if
  1. 1.
    $\hat{b}_{ij}<\hat{b}_{ik}+\hat{b}_{jk}$, for every distinct $i,j,k\in[3]$.
  2. 2.
    $\hat{b}_{12}+\hat{b}_{13}+\hat{b}_{23}$ is even.
Proof.
If for some tree $T$, $\hat{B}=\hat{D}_{T}$ is the ${\cal L}(T)$-distance matrix of $T$, then it is a routine exercise to check that $\hat{B}$ satisfies properties (1) and (2).
To prove the converse, let $\hat{B}$ be a dissimilarity matrix of order $3$:
${\hat{B}}=\left(\begin{array}[]{rrr}0&a&b\cr a&0&c\cr b&c&0\end{array}\right)$
Firstly, notice that $\min\{a,b,c\}\geq 2$, since otherwise if for example $\min\{a,b,c\}=a=1$, then, according to condition (1): $b<1+c$ and $c<1+b$, which means that $b=c$, and thus $a+b+c=1+2c$, contradicting condition (2).
Consider the tree $T$ of order $n=x+y+z+3$ with 3 leaves displayed in Figure 5 (left) and notice that if ${\hat{B}}={\hat{D}_{T}}$, then
$\left\{\begin{array}[]{rrrrrrrrrrrrrr}y+z=a\\ \\ x+z=b\\ \\ x+y=c\end{array}\right.\Leftrightarrow\left\{\begin{array}[]{rrrrrrrrrrrrrr}x=\frac{b+c-a}{2}\\ \\ y=\frac{a+c-b}{2}\\ \\ z=\frac{a+b-c}{2}\end{array}\right.$
Figure 5: Left: tree of order $n=x+y+z+3$ with 3 leaves. Right: tree of order $n=x+y+z+t+4$ with 4 leaves.
Clearly, $x$, $y$ and $z$ are strictly positive, since $\hat{B}$ satisfies property (1). Moreover, $x$, $y$ and $z$ are integers, since according to property (2), $a+b+c$ is an even integer, which means that integers $b+c-a$, $a+c-b$ and $a+b-c$ are also even. Hence, the ${\cal L}(T)$-distance matrix of $T$ is $\hat{B}$. ∎
And now, the case $4\times 4$.
Lemma 21.
Let $\hat{B}$ be an integer metric dissimilarity matrix or order 4 such that
  1. 1.
    $\hat{b}_{ij}<\hat{b}_{ik}+\hat{b}_{jk}$, for every distinct $i,j,k\in[4]$.
  2. 2.
    $\hat{b}_{ij}+\hat{b}_{ik}+\hat{b}_{jk}$ is even, for every distinct $i,j,k\in[4]$.
Then, $\hat{B}$ is the ${\cal L}(T)$-distance matrix of a tree $T$ and thus it is additive.
Proof.
Let $\hat{B}$ be a dissimilarity matrix of order $4$ satisfying properties (1) and (2):
${\hat{B}}=\left(\begin{array}[]{rrrr}0&a&b&\alpha\cr a&0&c&\beta\cr b&c&0&\gamma\cr\alpha&\beta&\gamma&0\end{array}\right)$
We may assume, without loss of generality, that
$a+\gamma\leq b+\beta\leq c+\alpha.$
Consider the tree $T$ of order $n=x+y+z+t+4$ with 4 leaves displayed in Figure 5 (right).
If ${\hat{B}}={\hat{D}_{T}}$, then according to Lemma 20:
$\left\{\begin{array}[]{rrrrrrrrrrrrrr}x=\frac{b+c-a}{2}\\ \\ y=\frac{a+c-b}{2}\\ \\ z=\frac{a+b-c}{2}\end{array}\right.$
Moreover:
$\left\{\begin{array}[]{rrrrrrrrrrrrrr}x=r+s\\ \\ \alpha=z+s+t\\ \\ \beta=y+s+t\\ \\ \gamma=r+t\end{array}\right.\Leftrightarrow\left\{\begin{array}[]{rrrrrrrrrrrrrr}r=\frac{x+z-\alpha+\gamma}{2}\\ \\ s=\frac{x-z+\alpha-\gamma}{2}\\ \\ t=\frac{-x-z+\alpha+\gamma}{2}\end{array}\right.\Leftrightarrow\left\{\begin{array}[]{rrrrrrrrrrrrrr}r=\frac{b-\alpha+\gamma}{2}\\ \\ s=\frac{c-a+\alpha-\gamma}{2}\\ \\ t=\frac{-b+\alpha+\gamma}{2}\end{array}\right.$
Clearly, $r$ and $t$ are strictly positive, since $\hat{B}$ satisfies property (1). Moreover, $r$ and $t$ are integers, since according to property (2), $b+\alpha+\gamma$ is an even integer, which means that integers $b-\alpha+\gamma$ and $-b+\alpha+\gamma$ are both also even.
From the previous assumption 3.2 we derive that $s$ is non-negative. In addition, $s$ must be an integer, since $c-a+\alpha-\gamma=(b+\alpha-\gamma)-(a+b-c)$ is, according to property (2), either even or 0.
Hence, we have shown that ${\hat{B}}={\hat{D}_{T}}$, which means, according to Proposition 12, that ${\hat{B}}$ is an additive matrix. ∎
Remark 22.
Notice that, in the above proof, $a+\gamma=b+\beta=c+\alpha.$ if and only if $s=0$. Hence, in this case $T$ is a spider (see Figure 5 (right)).
At this moment, we are in disposition to prove the general result that determines which matrices are the $\partial(T)$-distance matrix of a tree $T$.
Theorem 23.
Let $\hat{B}_{\kappa}$ be an integer metric dissimilarity matrix of order $\kappa\geq 3$. Then, $\hat{B}_{\kappa}$ is the ${\cal L}(T)$-distance matrix of a tree $T$ if and only if it is additive and
  1. 1.
    $\hat{b}_{ij}<\hat{b}_{ik}+\hat{b}_{jk}$, for every distinct $i,j,k\in[\kappa]$.
  2. 2.
    $\hat{b}_{ij}+\hat{b}_{ik}+\hat{b}_{jk}$ is even, for every distinct $i,j,k\in[\kappa]$.
Proof.
If for some tree $T$ with $\kappa$ leaves, $\hat{B}=\hat{D}_{T}$ is the distance matrix of the leaves of $T$, then it is a routine exercise to check that $\hat{B}$ is additive and satisfies properties (1) and (2).
To prove the converse, take an additive integer metric dissimilarity matrix $\hat{B}_{\kappa}$ of order $\kappa\geq 5$ satisfying properties (1) and (2). We distinguish cases.
Case 1.: For every 4-subset of indices $\{i,j,h,k\}\subseteq[\kappa]$,
$\hat{b}_{ij}+\hat{b}_{hk}=\hat{b}_{ih}+\hat{b}_{jk}=\hat{b}_{ik}+\hat{b}_{jh}.$
We claim that $\hat{B}_{\kappa}$ is the distance matrix $\hat{D}_{S_{\kappa}}$ of the set of leaves of a spider $S_{\kappa}$. To prove it, we proceed by induction on ${\kappa}\geq 4$. Notice that if $\kappa=4$, then according to Remark 22, our claim holds. Let $\kappa\geq 5$ and take the matrix $\hat{B}_{{\kappa}-1}$ obtained by deleting the last row and column of $\hat{B}_{\kappa}$.
By the inductive hypothesis, $\hat{B}_{{\kappa}-1}$ is the distance matrix $\hat{D}_{S_{\kappa-1}}$ of the set of leaves of a spider $S_{\kappa-1}$. Let ${\cal L}(S_{\kappa-1})=\{\ell_{1},\ldots,\ell_{\kappa-1}\}$ the set of leaves of $S_{\kappa-1}$, in such a way that for every $i,j\in[\kappa-1]$, $d(\ell_{i},\ell_{j})=[\hat{B}_{{\kappa}-1}]_{ij}$. Take the principal submatrix $\hat{B}_{4}$ of $\hat{B}_{{\kappa}}$ of order 4 obtained by considering the rows (and thus also the columns) in $\{1,2,3,\kappa\}$. According to Remark 22, there is a (unique) spider $S_{4}$ such that the distance matrix of its set of leaves $\{\ell_{1},\ell_{2},\ell_{3},\ell_{\kappa-1}\}$ is $\hat{B}_{4}$.
Let $S_{\kappa}$ the spider obtained by joining the spiders $S_{\kappa-1}$ and $S_{4}$. If ${\cal L}(S_{\kappa})=\{\ell_{1},\ldots,\ell_{\kappa-1},\ell_{\kappa}\}$, then for every $i,j\in[\kappa]$, $d(\ell_{i},\ell_{j})=\hat{b}_{ij}$, unless $14\leq i\leq\kappa-$ 1 and $j=\kappa$. Finally, if $14\leq i\leq\kappa-$ 1, then $d(\ell_{i},\ell_{\kappa})+d(\ell_{1},\ell_{2})=d(\ell_{1},\ell_{\kappa})+d(\ell_{i},\ell_{2})$. Thus, according to condition (3.2):
$d(\ell_{i},\ell_{\kappa})=d(\ell_{1},\ell_{\kappa})+d(\ell_{i},\ell_{2})-d(\ell_{1},\ell_{2})=\hat{b}_{1\kappa}+\hat{b}_{i2}-\hat{b}_{12}=\hat{b}_{i\kappa}.$
Case 2.: For some subset of indices $\{i,j,h,k\}\subseteq[\kappa]$,
$\hat{b}_{ij}+\hat{b}_{hk}<\hat{b}_{ih}+\hat{b}_{jk}=\hat{b}_{ik}+\hat{b}_{jh}.$
We proceed by induction on $\kappa\geq 3$. Cases $\kappa=3$ and $\kappa=4$ have been proved in Lemma 20 and Lemma 20, respectively. W.l.o.g., we assume that condition (3.2) holds for a 4-subset $\{1,j,h,\kappa\}$. Let $\hat{B}^{1}_{{\kappa}-1}$ and $\hat{B}^{\kappa}_{{\kappa}-1}$ be the matrices obtained by deleting the row (and thus also the column) 1 and $\kappa$ of $\hat{B}_{\kappa}$, respectively. Let $\hat{B}_{{\kappa}-2}$ be the matrix obtained by deleting rows (and thus also columns) 1 and $\kappa$ of $\hat{B}_{\kappa}$.
By the inductive hypothesis, $\hat{B}_{{\kappa}-2}$,$\hat{B}^{1}_{{\kappa}-1}$ and $\hat{B}^{\kappa}_{{\kappa}-1}$ are, respectively, the distance matrices of the set of leaves of three trees $T_{\kappa-2}$, $T^{1}_{\kappa-1}$ and $T^{\kappa}_{\kappa-1}$. Moreover, according to Theorem 19, $T_{\kappa-2}$ is a subtree of both $T^{1}_{\kappa-1}$ and $T^{\kappa}_{\kappa-1}$.
Let $T_{\kappa}$ the tree obtained by joining trees $T^{1}_{\kappa-1}$ and $T^{\kappa}_{\kappa-1}$. If ${\cal L}(T^{1}_{\kappa-1})=\{\ell_{2},\ldots,\ell_{\kappa}\}$, ${\cal L}(T^{\kappa}_{\kappa-1})=\{\ell_{1},\ldots,\ell_{\kappa-1}\}$ and ${\cal L}(T_{\kappa})=\{\ell_{1},\ldots,\ell_{\kappa-1},\ell_{\kappa}\}$, then for every $i,j\in[\kappa]$, $d(\ell_{i},\ell_{j})=\hat{b}_{ij}$, unless $i=1$ and $j=\kappa$. Finally, $d(\ell_{1},\ell_{\kappa})+d(\ell_{j},\ell_{h})=d(\ell_{1},\ell_{h})+d(\ell_{j},\ell_{\kappa})$. Thus, according to condition (3.2):
$d(\ell_{1},\ell_{\kappa})=d(\ell_{1},\ell_{h})+d(\ell_{j},\ell_{\kappa})-d(\ell_{j},\ell_{h})=\hat{b}_{1h}+\hat{b}_{j\kappa}-\hat{b}_{jh}=\hat{b}_{1\kappa},$
which means that $\hat{B}_{{\kappa}}$ is the distance matrix of $T_{\kappa}$. ∎
3.3 Reconstructing a tree from its leaf distances
To finalize this section, we show how to algorithmically reconstruct a tree $T$ from its ${\cal L}(T)$-distance matrix. To this end, it is enough to notice that the proof of Theorem 19 can be turned into an algorithm which runs in the worst case in $O(\kappa n)$ time.
Algorithm 1 Reconstructing-Tree
1:A matrix $\hat{D}_{T}$ of a certain tree.
2:A tree $T=(V,E)$.
3:Let $\kappa$ be the order of the matrix $\hat{D}_{T}$ and let $T$ be initially a set of $\kappa$ isolated vertices $\ell_{1},\ldots,\ell_{\kappa}$;
4:Join the vertices $\ell_{1}$ and $\ell_{2}$ by a path of the length determined in $\hat{D}_{T}$;
5:Label all the vertices $u\in V$ with $r(u|\{\ell_{1},\ell_{2}\})$, i.e., the distances from $u$ to $\{\ell_{1},\ell_{2}\}$;
6:for $k:=3$ to $\kappa$ do
7:     Compute $r(\ell_{k}|\{\ell_{1},\ldots,\ell_{k-1}\})$ as the distances from $\ell_{k}$ to $\{\ell_{1},\ldots,\ell_{k-1}\}$;
8:      Locate a vertex $u$ in $T$ and a positive integer $a$ such that $r(u|\{\ell_{1},\ldots,\ell_{k-1}\})+(a,\stackrel{{\scriptstyle k-1}}{{\ldots}}a)=r(\ell_{k}|\{\ell_{1},\ldots,\ell_{k-1}\})$;
9:     Add to $T$ a path of length $a$ joining $u$ and $\ell_{k}$;
10:     Relabel all the vertices in $T$ with their distances to $\{\ell_{1},\ldots,\ell_{k}\}$;
11:end for
12:return $T$.
Corollary 24.
The Algorithm 1 runs in time $O(\kappa n)$.
Proof.
It is straightforward to check that the step dominating the computation is 8, and that step is repeated $O(\kappa n)$ times. ∎
4 Reconstructing unicyclic graphs from the boundary
In this section, we move from trees to unicyclic graphs, i.e., those graphs containing a unique cycle. It is divided into two subsections: one devoted to the procedure for knowing whether a matrix is the distance matrix of a unicyclic graph or not. The second one is dedicated to the process of reconstructing a unicyclic graph with $g\geq 5$ from its $\partial(G)$-distance matrix.
4.1 The distance matrix of a unicyclic graph
In this subsection, our goal is to recognize whether a matrix is the distance matrix of a unicyclic graph. Unlike in the previous section, the procedure to check that condition is algorithmic. At each step, one can delete a leaf from the graph, which is easily recognizable since it has only a one in its row and column. When there are no leaves, the resulting matrix should be one of a cycle (see Figure 6).
Figure 6: Procedure to recognize distance matrices of unicyclic graphs.
Theorem 25.
A graph $G$ is unicyclic if and only the algorithm described above returns True.
Proof.
Let $D_{G}$ be the distance matrix of a graph $G$. Then, a row and a column with a unique one corresponds with a leaf in the graph $G$, and if we delete that row and column then the new matrix $D^{\prime}$ is the distance matrix of the graph $G^{\prime}$ obtained by deleting that leaf in $G$ (see Figure 6). Hence, if the final matrix of the above procedure is the distance matrix of a cycle, then it only remains to rebuild the graph to obtain a unicyclic graph. ∎
4.2 Reconstructing a unicyclic graph from its boundary distances
In this subsection, it is described the process of reconstructing a unicyclic graph from the distance matrix of its boundary. The idea of the algorithm is the following: from the matrix $\hat{D}_{G}$ we recognize the different branching trees of the graph $G$. Then, we can apply Algorithm 1 of trees to the reconstruction of those subtrees. On the other hand, we simplify $\hat{D}_{G}$ to contain all the vertices of the cycle and reconstruct it. Only remains to join the cycle with the subtrees to have again the unicyclic graph.
According to Proposition 6, $\partial(G)={\cal L}(G)\cup{\cal U}(G)$. The next result gives us a procedure to distinguish the vertices of $\partial(G)$ with just the information provided by the matrix $\hat{D}_{G}$.
Lemma 26.
Let $G$ be a unicyclic graph with $g\geq 4$. Given the matrix $\hat{D}_{G}$, it is possible to distinguish the vertices in ${\cal L}(G)$ from the ones in ${\cal U}(G)$.
Proof.
First of all, if there is a one in a certain position of $\hat{D}_{G}$, then the two vertices of $\partial(G)$ at that position are adjacent, and so the vertices belong to ${\cal U}(G)$. Let $u\in{\cal U}(G)$ be one vertex not being in the above situation. Then, it has two neighbours with a branching tree having each of them at least one leaf $v_{1}$ and $v_{2}$. Remember that $v_{1},v_{2}\in{\cal L}(G)$. Since $g\geq 4$, the shortest path between $v_{1}$ and $v_{2}$ goes through $u$ and that can be detected in $\hat{D}_{G}$ by checking whether $d(v_{1},v_{2})=d(v_{1},u)+d(u,v_{2})$. Reciprocally, if three vertices $v_{1},v_{2},u\in\partial(G)$ verify that $d(v_{1},v_{2})=d(v_{1},u)+d(u,v_{2})$, and again since $g\geq 4$, then $u$ is in the shortest path joining $v_{1}$ and $v_{2}$ and cannot be a leaf, and thus $u\in{\cal U}(G)$.
Once the vertices of ${\cal U}(G)$ have been located, by Proposition 6, ${\cal L}(G)=\partial(G)\setminus{\cal U}(G)$. ∎
In the process described above, we will need the information of the branching vertex of a branching tree which originally is never a boundary vertex. However, the next result shows how we can obtain that information from $\hat{D}_{G}$.
Lemma 27.
Let $G$ be a unicyclic graph and let $u\in{\cal L}(G)$. It is possible to determine the length of the path ending at $u$ only with the information provided in $\hat{D}_{G}$.
Proof.
Suppose that $u$ hangs out from the vertex $x$. Since $u\in{\cal L}(G)$, clearly $x\notin{\cal U}(G)$. We can delete $u$ from $G$, which in terms of the matrix $\hat{D}_{G}$ it means that the row and column corresponding to $u$ are reduced by one except the zero in the position $(u,u)$. Then, we can check if $x\in{\cal U}(G)$ using the procedure described in Lemma 26. If not, we continue deleting the vertices of the path between $u$ and $x$ until $x\in{\cal U}(G)$. The number of deleted vertices is the length of the path. ∎
The next lemma is key for the procedure we are looking for. It provides a criteria for separating the leaves of different branching trees. However, this criteria only work for unicyclic graphs of girth at least $5$. In obtaining this result, we use again the concept of doubly resolving sets.
Lemma 28.
Let $G$ a unicyclic graph with girth $g\geq 5$. Let $u\in{\cal L}(G)$ and let $S_{u}\subseteq V(G)$ be such that $u\in S_{u}$. Then, $S_{u}$ is the set of leaves of $T_{u}$ if and only if for all $w\in S_{u}$ there exists an integer $a_{w}$ such that $r(w|\partial(G)\setminus S_{u})=(a_{w},\stackrel{{\scriptstyle\kappa-|S_{u}|}}{{\ldots}},a_{w})+r(u|\partial(G)\setminus S_{u})$.
Proof.
Consider a certain leaf $u$ and the set $S_{u}$ of leaves in the same branching tree having $x\in V(G)$ as its branching-vertex, and $w\in S_{u}$. Then, we have that the distance between $u$ and $y\in\partial(G)\setminus S_{u}$ is $d(u,y)=d(u,x)+d(x,y)$ and since $w$ is in the same branching tree as $u$, the same occurs for $w$, i.e., $d(w,y)=d(w,x)+d(x,y)$. So, $d(w,y)-d(u,y)=d(w,x)-d(u,x)=a_{w}$ and therefore $d(w,y)=a_{w}+d(u,y)$. Note that $a_{w}$ does not depend of $y$ so it is the same integer for all the vertices in $\partial(G)\setminus S_{u}$.
Reciprocally, let $u$ be a leaf on $V(G)$ and $S_{u}$ is the set of leaves in the same branching tree as $u$. Let $w\in V(G)$ another leaf such that $r(w|\partial(G)\setminus S_{u})=(a_{w},\stackrel{{\scriptstyle\kappa-|S_{u}|}}{{\ldots}},a_{w})+r(u|\partial(G)\setminus S_{u})$. However, on the contrary, assume that $w\notin S_{u}$ which means that $w$ belongs to a different branching tree as $u$ with a different branching vertex $z$. There are three possibilities here: either $x$ and $z$ are adjacent or $d(x,z)=2$ or $d(x,z)>2$.
First, let us suppose that $x$ and $z$ are adjacent. Since $g\geq 5$, there exist two vertices in the cycle $x^{\prime}$ and $z^{\prime}$, respectively adjacent to $x$ and $z$. For the sake of simplicity, we might assume that $x^{\prime},z^{\prime}$ are not branching vertices. Otherwise, we can make the same reasoning with a leaf in each branching tree. Hence,
$d(u,x^{\prime})=d(u,x)+1;\qquad$ $d(w,x^{\prime})=d(w,z)+2$
$d(u,z^{\prime})=d(u,x)+2;\qquad$ $d(w,z^{\prime})=d(w,z)+1$
According to the above, the value of $a_{w}$ would be $d(w,x^{\prime})-d(u,x^{\prime})=d(w,z)-d(u,x)+1$ and $d(w,z^{\prime})-d(u,z^{\prime})=d(w,z)-d(u,x)-1$, which is clearly impossible.
On the other hand, let us suppose that $d(x,z)>2$, i.e., that there exist at least two intermediate vertices $y$ and $y^{\prime}$ in the shortest path between $x$ and $z$ being $y$ closer to $x$ than to $z$. Again, for simplicity, we can assume that neither $y$ nor $y^{\prime}$ are branching vertices since otherwise we can reasoning with a leaf hanging out from each of them. Then, $d(w,y)-d(u,y)=d(w,z)+d(z,y^{\prime})+d(y^{\prime},y)-d(u,x)-d(x,y)$ and $d(w,y^{\prime})-d(u,y^{\prime})=d(w,z)+d(z,y^{\prime})-d(u,x)-d(x,y)-d(y,y^{\prime})$ which cannot be equal unless $y=y^{\prime}$ and $d(x,z)=2$, which is not the case.
Finally, suppose that $d(x,z)=2$, i.e., that there exists only one intermediate vertex $y$ in the shortest path between $x$ and $z$. For the same reasons as above, we may assume that $y$ is not a branching vertex. Since $g\geq 5$, there is another vertex in the cycle $x^{\prime}$ adjacent to $x$. Then, $d(w,y)-d(u,y)=d(w,z)-d(u,x)$ and at the same time $d(w,x^{\prime})-d(u,x^{\prime})=d(w,z)+d(z,x^{\prime})-d(u,x)-1$ and those two quantities are only equal if $d(z,x^{\prime})=1$, which is not possible since $g\geq 5$. Thus, in any case we have reached a contradiction and so necessarily $w\in S_{u}$. ∎
Note that the hypothesis $g\geq 5$ is necessary as otherwise we could have the graph of Figure 6 that does not verify Lemma 28.
At this point, we are able to give a pseudocode description of the algorithm. Note the uses of the previous results to obtain the final output. Finally, Theorem 29 establishes the correctness and time complexity of Algorithm 2.
Algorithm 2 Reconstructing-Unicyclic
$\hat{D}_{G}$ of a unicyclic graph $G$.
The graph $G$.
Classify the vertices of $\hat{D}_{G}$ according to Lemma 26;
Separate the vertices of each branching tree $T_{u}$ according to Lemma 28;
Add the branching vertex $u$ and find the distance of it to the rest of vertices of its branching tree with Lemma 27;
Store the distances between vertices of $T_{u}$ in a submatrix $\hat{D}_{u}$;
Delete all the boundary leaves from $\hat{D}_{G}$ and add the branching vertices to it;
if the remaining $\hat{D}_{G}$ is not the matrix of distances of a cycle then
     return “The matrix does not correspond to a unicyclic graph”;
else
     Let $G$ be the cycle that corresponds with the matrix;
     for each branching vertex $u$ do
         Apply Algorithm 1 to $\hat{D}_{u}$ to obtain the subtree $T_{u}$;
         Add the subtree $T_{u}$ to $G$;
     end for
end if
return $G$.
Theorem 29.
Beginning with the boundary distance matrix $\hat{D}_{G}$ of a unicyclic graph $G$ with $g\geq 5$, Algorithm 2 obtains an isomorphic graph to $G$ in $O(n^{2})$.
Proof.
Once the boundary vertices that are leaves have been marked as shown in Lemma 26, the algorithm gathers the leaves of the same subtree $T_{u}$ using the characterization provided in Lemma 28. Then, a new submatrix is constructed for the leaves of $T_{u}$ including $u$ which originally is not a boundary vertex. Now, we are in disposition to use Algorithm 1 to reconstruct the subtree $T_{u}$.
At the same time, the leaves of each branching tree $T_{u}$ are deleted from $\hat{D}_{G}$ and the branching vertices as $u$ are added. The resulting matrix should be the distance matrix of a cycle. It only remains to construct that cycle add joined it with the different subtrees $T_{u}$.
The complexity is given by several steps that run in time $O(n^{2})$. ∎
5 Conclusions and Further work
In [22], it was firstly implicitly mentioned that a resolving set $S$ of a graph $G$ is strong resolving if and only if the distance matrix $D_{S,V}$ uniquely determines the graph $G$ (see Theorem 9). On the other hand, in [21] it was proved that the boundary $\partial(G)$ of every graph $G$ is a strong resolving set (see Proposition 4).
Mainly having in mind this pair of results, we have presented in Section 2 the following conjecture.
Conjecture 30.
Every graph belongs to ${\cal H}_{2}$.
In Sections 3 and 4, we have proved that every pseudotree belongs to ${\cal H}_{2}$, and we have also provided algorithms to recognize both trees and unicyclic graphs.
In addition, in Section 3, we have been able to characterize, for trees, both the distance matrix $D_{T}$ and the $\partial(T)$-distance matrix $\hat{D}_{T}$ (see Theorems 18, 19 and 23).
We conclude with a list of suggested open problems.
Open Problem 1:
Characterizing distance matrices of unicyclic graphs in a similar way as it has been done for trees in this work.
Open Problem 2:
Solving the case of reconstructing unicyclic graphs with girth at most 4.
Open Problem 3:
Checking whether every block graph belongs to ${\cal H}_{1}$, or at least to ${\cal H}_{2}$.
Open Problem 4:
Checking whether every cactus graph belongs to ${\cal H}_{1}$, or at least to ${\cal H}_{2}$.
Open Problem 5:
Checking whether every block-cactus graph belongs to ${\cal H}_{1}$, or at least to ${\cal H}_{2}$.
Open Problem 6:
Checking whether every split graph belongs to ${\cal H}_{1}$, or at least to ${\cal H}_{2}$.
Open Problem 7:
Checking whether every Ptolemaic graph belongs to ${\cal H}_{1}$, or at least to ${\cal H}_{2}$.
Open Problem 8:
Checking whether every graph of diameter 2 belongs to ${\cal H}_{2}$, and characterizing the set of graphs of diameter 2 belonging to ${\cal H}_{1}$.
Open Problem 9:
Checking whether every graph of diameter 3 belongs to ${\cal H}_{2}$, and characterizing the set of graphs of diameter 3 belonging to ${\cal H}_{1}$.